OpenStudy (anonymous):

A jogger and a cyclist set out at 9 A.M. from the same point headed in the same direction. The average speed of the cyclist is four times the average speed of the jogger. In 2 h, the cyclist is 33 mi ahead of the jogger. How far did the cyclist ride?

OpenStudy (anonymous):

I am stuck ! :(

OpenStudy (anonymous):

Let x = the average speed of the jogger. Let y = the average speed of the cyclist. y = 4x 2y = 2x + 33 Substitute 4x for y and solve for x 2(4x) = 2x + 33 8x = 2x + 33 6x = 33 x = 33/6 = 5.5 <-- this is the average speed of the jogger Plug 5.5 in for x in the equation y = 4x y = 4(5.5) = 22 <-- this is the average speed of the cyclist Now plug both values into the equation 2y = 2x + 33 2(22) = 2(5.5) +33 44 = 11 + 33 44 = 44 <-- this is the total distance traveled by the cyclist